OpenStudy (anonymous):

The length of the hypotenuse of a right triangle is 137 units. The length of one leg of the triangle is 105 units. Gina wrote the following step to find the length of the unknown leg: Length of the unknown leg = 1372 − 1052 = 18,769 − 11,025 = 7,744 units Which statement best explains whether Gina's step is correct or incorrect?

OpenStudy (anonymous):

It is incorrect because the length of the unknown side is the square root of 29,794. It is incorrect because the length of the unknown side is the square root of 7,744. It is correct because the length of the unknown side is the difference of the lengths of the sides. It is correct because the length of the unknown side is the difference of the squares of the sides.

OpenStudy (anonymous):

@10115658

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

x^2 = 18769 - 11025

OpenStudy (anonymous):

what should x be?

OpenStudy (anonymous):

b=88

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

So whats the answer

OpenStudy (anonymous):

whats 8x8?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Did you use that site to solve or did you do it algebraically?

OpenStudy (anonymous):

x is 7744?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

so we have x^2 = 7744

OpenStudy (anonymous):

but 88 x 88 is 7744

OpenStudy (anonymous):

what is the last step

OpenStudy (anonymous):

The last step is to take the square root of both sides

OpenStudy (anonymous):



OpenStudy (anonymous):